OpenStudy (anonymous):

a point whose abscissa and ordinate are equalis square root of 10 units from (-1,-3). find the points

OpenStudy (anonymous):

\[\sqrt{(x+1)^{2}+(y+3)^{2}}=\sqrt{10}\]

OpenStudy (anonymous):

do u have any options or u need to solve it completely

OpenStudy (anonymous):

?

OpenStudy (anonymous):

ahmmm.. i justwant to know HOW to solve that question :))

OpenStudy (anonymous):

see as i have the equation...........one way is you can arbitrary values of x and y and check when the left hand side is equal to square root of 10

OpenStudy (anonymous):

**arbitrarily substitute

OpenStudy (anonymous):

like it can be (2,-2)

OpenStudy (anonymous):

errr.. the books says it was 0,0 and -4, -4

OpenStudy (anonymous):

ya even both these points satisfy the given equation

OpenStudy (anonymous):

i just gave u an example ......

OpenStudy (anonymous):

but the condition is both the abscissa and the ordinate are equal.. :))

OpenStudy (anonymous):

ok ...............got in this equation put x=y .............you will get a quadratic in x nad then solve to get two values of x and two values of y

OpenStudy (anonymous):

(x+1)^2+(x+3)^2=10 \[2x ^{2}+8x+10=10\] \[2x ^{2}+8x=0\] 2x(x+4)=0 x=0 or x=-4 since x=y, thus,(0,0) and (-4,-4)

OpenStudy (anonymous):

hmmmmmm... i will consider that.. by the way, thanks! i now have a direction on how to solve this stuff.. i will inform you if i solve that stuff :))

OpenStudy (anonymous):

see you had two equations:: \[\sqrt{(x+1)^{2}+(y+3)^{2}}=\sqrt{10}\] and x=y

OpenStudy (anonymous):

wiow! :0

OpenStudy (anonymous):

wow! now i know wheres my error .. thank you so much! you really help me in this stuff :D

OpenStudy (anonymous):

ur welcome......